OpenStudy (anonymous):

2/3-1/4 i got 23/32 this is how 2/3-1/4=2/3*12/12=24/36-1/4=23/32... that is not right so what did i do wrong

OpenStudy (anonymous):

why didnt you convert the 1/4 also to something/36? you cannot simply subtract fractions if they dont have the same denominator.

OpenStudy (anonymous):

well i thought that by multiplying by the LCM which is 12 i was doing well atlest up til that point was I right?

OpenStudy (anonymous):

oh i should have used 3

OpenStudy (anonymous):

instead of 12

OpenStudy (anonymous):

yes, but you need to multiply BOTH fractions by the LCM, not just the one you like.

OpenStudy (anonymous):

the LCM of 3 and 4 is 12 so 2/3 can be expressed as \[\frac 2 3 * \frac 4 4 = \frac 8 {12}\]

OpenStudy (anonymous):

similarly 1/4 can be expressed as \[\frac 1 4 \times \frac 3 3 = \frac 3{12}\]

OpenStudy (anonymous):

now that both denominators are the same, you can subtract the numerators

OpenStudy (anonymous):

what do you get when you subtract one-third from two-thirds?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

how did you work that out? did you do 2/3 - 1/3 = 2-1 / 3-3 = 2/0 = 2?

OpenStudy (anonymous):

subtracted across was that wrong

OpenStudy (anonymous):

yes first of all it doesnt make sense. two thirds is 2/3 which is 0.66666666 one third is 1/3 or 0.3333333333 2/3 - 1/3 = 0.6666666-0.33333333 = 0.3333333, not 2